OpenStudy (anonymous):

a building casts a shadow 230 meters long at the same time a pole 2 meters high casts a shadow 10 meters long what is the height of the building?

OpenStudy (anonymous):

use ratio and proportion height of the pole/ shadow length = height of the building/ shadow length 2/10=x/230 10x= 2(230) x= 46 meters. .

OpenStudy (anonymous):

the height of the building is 46 meters. .

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

glad to help. .